\begin{thebibliography}{1}



\bibitem{LMD} L.M. DeAlba,
\newblock  Acyclic and unicyclic graphs whose minimum skew rank is equal to the minimum skew rank of
a diamtrical path,
\newblock Preperint, arXiv:1107.2170v1.

\bibitem{DKT} L.M. DeAlba, E. Kerzner, S. Tucker,
\newblock  A note on the minimum skew rank of powers of
paths,
\newblock Preprint. arXiv:1107.2450v1.

\bibitem{De}  L. DeLoss.
\newblock  Results on minimum skew rank of matrices described by a
graph,
\newblock Masters Thesis, Iowa State University,  May 2009.

\bibitem{FH} S. Fallat, L. Hogben,
\newblock  The minimum rank of symmetric matrices described by a graph: A
survey.
\newblock   Linear Algebra Appl. 426 (2007) 558--582.


\bibitem{AB} IMA-ISU research group on minimum rank
(M. Allion, E. Bodine, L.M. DeAlba, J. Debnath, L. DeLoss, C.
Garnett, J. Grout, L. Hogben, B. Im, H. Kim, R. Nair, O. Pryporova,
K. Savage, B. Shader, A.W. Wehe),
\newblock  Minimum rank of skew-symmetric
matrices described by a graph,
\newblock   Linear Algebra Appl. 432 (2010)
2457--2472.

\bibitem{Ro} D.J. Rose,
\newblock  On simple characterizations of $k$-trees,
\newblock   Discrete Math.  7 (1974) 317--322.

\bibitem {Sh} D.R. Shier,
\newblock  Some aspects of perfect elimination orderings in chordal graphs,
\newblock  Discrete Appl. Math. 7 (1984) 325--331.

\bibitem{We}
D.B. West,
\newblock   Introduction to Graph Theory,
\newblock  Prentice Hall, New Jersey,
2001.

\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{amsmath,amsfonts,amssymb,amscd}
\newtheorem {Lemma}{Lemma}[section]
\newtheorem {Theorem}{Theorem}[section]

\newtheorem {Corollary}{Corollary}[section]
\newenvironment {proof} {\noindent {\bf Proof.}}{\hspace*{\fill}$\Box$\par\vspace{4mm}}

\begin{document}

\title{
A note on the minimum skew rank of  a graph}
\author{
Yanna Wang and Bo Zhou\\
Department of Mathematics, South China Normal
University,\\ Guangzhou 510631, P. R. China\\
e-mail: zhoubo@scnu.edu.cn}
\date{}
\maketitle

\begin{abstract}
The minimum skew rank $mr^{-}(\mathbb{F},G)$ of a graph $G$ over a
field $\mathbb{F}$ is the smallest possible rank among all skew
symmetric matrices over $\mathbb{F}$, whose ($i$,$j$)-entry (for
$i\neq j$) is nonzero whenever $ij$ is an edge in $G$ and is zero
otherwise. We give some new properties of the minimum skew rank of a graph, including
a characterization of the graphs $G$ with cut vertices over the
infinite field $\mathbb{F}$ such that $mr^{-}(\mathbb{F},G)=4$,
determination of the minimum skew rank of $k$-paths over a field $\mathbb{F}$, and an extending of an existing result to show that
$mr^{-}(\mathbb{F},G)=2match(G)=MR^{-}(\mathbb{F},G)$ for a connected graph $G$ with no
even cycles and a field $\mathbb{F}$, where $match(G)$ is the
matching number of $G$, and $MR^{-}(\mathbb{F},G)$ is the largest possible rank among all skew
symmetric matrices over $\mathbb{F}$.
\\ \\
{\bf Key words:}
Minimum skew rank, Skew-symmetric  matrix, $k$-tree, $k$-path,
Zero forcing number, Perfect matching\\
\\
{\bf AMS Classifications:}
05C50, 15A03
\end{abstract}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Introduction}

We consider only simple graphs. Let $G$ be a graph with vertex set
$V_G$ and edge set $E_G$. Let $\mathbb{F}$ be a field. We adopt the
notation and terminology from \cite{AB} and \cite{We}.

An $n\times n$ matrix $A$ over $\mathbb{F}$ is skew-symmetric
(respectively, symmetric) if $A^T=-A$ (respectively, $A^T = A$),
where $A^T$ denotes the transpose of $A$.

For an $n\times n$ symmetric or skew-symmetric  matrix $A$, the
graph of $A$, denoted $G(A)$, is the graph with vertex set
$\{v_{1},v_{2},\dots, v_{n}\}$ and edge set $\{v_{i}v_{j}:
a_{ij}\neq 0, 1\leq i <j \leq n\}$. %The diagonal is ignored in
%determining $G(A)$ if $A$ is  symmetric, while the diagonal must be
%$0$ if $A$ is skew-symmetric.

The classic minimum rank problem involving symmetric matrices has
been studied extensively, see, e.g.,~\cite{FH}.

The minimum skew rank problem involves skew symmetric matrices  and
its study began recently in \cite{AB}. 
%see \cite {LMD,DKT} for more results. 
If the characteristic of $\mathbb{F}$ is $2$, then a
skew-symmetric matrix over $\mathbb{F}$ is also symmetric. %and may
%have nonzero diagonal entries.
Thus it is assumed throughout this
paper that the characteristic of $\mathbb{F}$ is not $2$.

For a field $\mathbb{F}$ and a graph $G$, let
$S^{-}(\mathbb{F},G)=\{A\in \mathbb{F}^{n\times n}: A^{T}=-A,
G(A)=G\}$ be the set of skew-symmetric matrices over $\mathbb{F}$
described by $G$. The minimum skew rank of $G$ over $\mathbb{F}$ is
defined as
\[
mr^{-}(\mathbb{F},G) = \min\{\text{rank}(A): A\in
S^{-}(\mathbb{F},G)\}.
\] The
corresponding maximum skew nullity of $G$ is defined as
\[ M^{-}(\mathbb{F},G)
= \max\{\text{nullity}(A): A\in S^{-}(\mathbb{F},G)\}.
\]
Obviously, $mr^{-}(\mathbb{F},G)+M^{-}(\mathbb{F},G)=|V_G|$.

Let $K_n$ be the complete graph with $n$ vertices, and $K_{n_1,n_2,
\dots, n_t}$ the complete $t$-partite graph with $n_i$ vertices in
the $i$th partite sets for $i=1, 2, \dots, t$.

Note that the rank of a skew-symmetric matrix over $\mathbb{F}$ is
always even. Thus $mr^{-}(\mathbb{F},G)$ is even for any field
$\mathbb{F}$ and any graph $G$. As observed in \cite{AB},
$mr^{-}(\mathbb{F},G)=0$ if and only if $G$ is an empty graph, and
if $\mathbb{F}$ is infinite and $G$ is a connected graph with at
least two vertices, then $mr^{-}(\mathbb{F},G)=2$ if and only if
 $G$ is a complete multipartite graph $K_{n_{1},n_{2},\ldots , n_{t}}$ for some $t\geq 2$,
$n_{i}\geq 1$ for $i=1,\ldots, t$. The authors \cite{AB} posed an
open question (Question 5.2) to characterize the graphs $G$ such
that $mr^{-}(\mathbb{F},G)=4$. We characterize the graphs $G$ with cut vertices over
the infinite field $\mathbb{F}$ such that $mr^{-}(\mathbb{F},G)=4$.

The class of $k$-trees is defined recursively as follows \cite{Ro}:
(i) The complete graph $K_{k+1}$ is a $k$-tree; (ii) A $k$-tree $G$
with $n+1$ vertices ($n\ge k+1$) can be constructed from a $k$-tree
$H$ on $n$ vertices by adding a vertex adjacent to all vertices of a
$k$-clique of $H$.
%
%A $k$-tree is a graph that can be built up from a complete graph $K_{k+1}$ on $k+1$ vertices by
%adding one vertex at a time adjacent to exactly the vertices in an
%existing $K_{k}$.
%A graph is a partial $k$-tree if it is a subgraph of a $k$-tree.
 A $k$-path is a $k$-tree which
is either $K_{k+1}$ or has exactly two vertices of degree $k$. We determine
the minimum skew rank of $k$-paths over a field $\mathbb{F}$. The $k$-th power $G^{k}$ of a graph $G$ is the graph
whose vertex set is $V_G$, two distinct vertices being adjacent in
$G^{k}$ if and only if their distance in $G$ is at most $k$. Let
$P_n=v_1v_2\dots v_n$ be the path on $n$ vertices.
If 
$k\le n-1$, then  $P_{n}^{k}$ is a
$k$-path (see below). As a corollary, we obtain the minimum skew rank
of the $k$-th power of a path  over the real field $\mathbb{R}$, which was already given in \cite{DKT}.

%
%For positive integer $k$, the $k$-th power of a graph $G$, denoted
%by $G^{k}$, is the graph with vertex set $V_G$ such that $uv\in
%E_{G^{k}}$  if and only if distance between $u$ and $v$ in $G$ is at
%most $k$.
% DeAlba et al.~\cite{DKT} determined

The maximum skew rank $MR^{-}(\mathbb{F},G)$ of a graph $G$ over a
field $\mathbb{F}$ is defined as
\[
MR^{-}(\mathbb{F},G) = \max\{\text{rank}(A): A\in
S^{-}(\mathbb{F},G)\}.
\]
Let $match(G)$ be the matching number of  $G$. It was shown
in \cite{AB} that
$mr^{-}(\mathbb{F},G)=2match(G)=MR^{-}(\mathbb{F},G)$ for a tree (a connected graph with no cycles) $G$
and a field $\mathbb{F}$. We extend this by showing that the above conclusion holds also for 
a
connected graph $G$ with no even cycles.

%
%In this paper, we characterize the graphs $G$ with cut vertices over
%the infinite field $\mathbb{F}$ such that $mr^{-}(\mathbb{F},G)=4$, determine
%the minimum skew rank of $k$-paths over a field $\mathbb{F}$, and show that
%$mr^{-}(\mathbb{F},G)=2match(G)=MR^{-}(\mathbb{F},G)$ for a
%connected graph $G$ with no even cycles and a field $\mathbb{F}$.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Preliminaries}

Let $G$ be a graph. For $v\in V_G$, $G-v$ denotes the graph obtained
from $G$ by deleting vertex $v$ (and all edges incident with $v$).
 For $X\subseteq V_G$, $G[X]$ denotes the
subgraph of $G$ induced by vertices in $X$.

We give some lemmas that we will use in our proof.

\begin{Lemma} \label{lm2.5}
\cite{AB} Let $G$ be a connected graph with at least two vertices
and let $\mathbb{F}$ be an infinite field.
 Then $mr^{-}(\mathbb{F},G)=2$ if and only if
 $G$ is a complete multipartite graph.% = K_{n_{1},n_{2},\ldots , n_{t}}$ for some $t\geq 2$,
%$n_{i}\geq 1$, $i=1,\ldots, t$.
\end{Lemma}

For a field $\mathbb{F}$  and a graph $G$ with $v\in V_G$, let
$r_{v}^{-}(\mathbb{F},
G)=mr^{-}(\mathbb{F},G)-mr^{-}(\mathbb{F},G-v)$.

The union of graphs $G_i$, $i=1,2, \dots,h$, denoted by
$\cup_{i=1}^{h}G_i$, is the graph with vertex set
$\cup_{i=1}^{h}V_{G_i}$ and edge set $\cup_{i=1}^{h}E_{G_i}$.

\begin{Lemma} \cite{AB,De}
 \label{lm2.7}
 Let $G$ be a graph with cut vertex $v$ and  $\mathbb{F}$  a field, where
$G=\cup_{i=1}^{h}G_i$ and $\cap_{i=1}^{h}V_{G_i}=\{v\}$. Then
$mr^{-}(\mathbb{F},G)=\sum_{i=1}^{h}mr^{-}(\mathbb{F},G_{i}-v)+ \min
\{\sum_{i=1}^{h}r_{v}^{-}(\mathbb{F},G_{i}),2\}$.
\end{Lemma}

\begin{Lemma} \label{lm2.1} \cite{AB}
Let $G$ be a graph and let $\mathbb{F}$ be an infinite field. If $G
=G_1\cup G_2$, then $mr^{-}(\mathbb{F},G) \leq
mr^{-}(\mathbb{F},G_1)+mr^{-}(\mathbb{F},G_2)$.
\end{Lemma}

Let $G$ be a graph. A subset $Z\subset V_G$ defines an initial
coloring by coloring all vertices in $Z$ black and all the vertices
outside $Z$ white. The color change rule says: If a black vertex $u$
has exactly one white neighbor $v$, then change the color of $v$ to
black. In this case we write $u\rightarrow v$. The derived set of an
initial coloring $Z$ is the set of vertices colored black until no
more changes are possible. A zero forcing set is a subset $Z\subset
V_G$ such that the derived set of $Z$ is $V_G$. The zero forcing
number of $G$, denoted by $Z(G)$, is the minimum size of a zero
forcing set of $G$.

\begin{Lemma} \label{lm2.4}
\cite{AB} Let $G$ be a graph  and
$\mathbb{F}$  a field. Then  $M^{-}(\mathbb{F},G)\leq Z(G)$.
\end{Lemma}

\begin{Lemma}\label{lm2.8} \cite{AB}
Let $G$ be a graph  and  $\mathbb{F}$  a
field. Then  $MR^{-}(\mathbb{F},G)=2match(G)$.
\end{Lemma}

\begin{Lemma}\label{lm2.9} \cite{AB} Let $G$ be a graph  and  $\mathbb{F}$  a
field. If $H$ is an induced subgraph of $G$,
$mr^{-}(\mathbb{F},H)\leq $ $ mr^{-}(\mathbb{F},G)$.
\end{Lemma}

\begin{Lemma} \label{lm2.6}
\cite{AB} Let $G$ be a graph with a unique perfect matching and
$\mathbb{F}$ a field. Then $mr^{-}(\mathbb{F},G) =|V_G|$.
\end{Lemma}

\section{Results}
%The minimum skew rank of $k$-path and cycle with chords}

we first give a partial result to the question in \cite{AB} to
characterize graphs with $mr^{-}(\mathbb{F},G)=4$. We consider
graphs with cut vertices.

\begin{Theorem} Let $G$
be a graph with cut vertex $v$ and  $\mathbb{F}$ an infinite field.
Then $mr^{-}(\mathbb{F},G)=4$ if and only if one of the following
conditions holds:\\
$(i)$ $G =G_1\cup G_2$ and $V_{G_1}\cap V_{G_2}=\{v\}$, where $G_1$,
$G_2$ are complete multipartite graphs such that $G_{1}-v$,
$G_{2}-v$ are
nonempty, and\\
$(ii)$ $G-v$ consists of a nonempty complete multipartite component
and isolated vertices.
\end{Theorem}

\begin{proof}
Suppose first that  (i) holds. Note that $G_{i}-v$ is still a
complete multipartite graph for $i=1,2$. By Lemma~\ref{lm2.5},
$mr^{-}(\mathbb{F},G_{1})=mr^{-}(\mathbb{F},G_{2})=mr^{-}(\mathbb{F},G_{1}-v)=mr^{-}(\mathbb{F},G_{2}-v)=2$.
Then $r_{v}^{-}(\mathbb{F},G_{1})+r_{v}^{-}(\mathbb{F},G_{2})=0$.
Thus by Lemma~\ref{lm2.7},
$mr^{-}(\mathbb{F},G)=mr^{-}(\mathbb{F},G_{1}-v)+mr^{-}(\mathbb{F},G_{2}-v)+\min\{0,2\}=4$.

Now suppose that  (ii) holds. Let $W$ be the unique complete
multipartite component, and $a$ the number of isolated vertices in
$G-v$. By Lemma~\ref{lm2.5}, $mr^{-}(\mathbb{F},W)=2$. Note that
$r_v^{-}(\mathbb{F},K_{2})=2$. Then by Lemma~\ref{lm2.7},
$mr^{-}(\mathbb{F},G)=mr^{-}(\mathbb{F},W)+a\cdot
mr^{-}(\mathbb{F},K_{1})+2=4$.

Conversely, suppose that  $mr^{-}(\mathbb{F},G)=4$. Let $p$ be the
number of nonempty complete multipartite components, and $q$ the
number of isolated vertices in $G-v$. Let $m$ be the number of the
remaining components.

Note that the minimum skew rank of a graph that is neither a
complete multipartite graph nor an empty graph is larger than $4$.

\noindent {\bf Case 1.} $q=0$. By Lemma~\ref{lm2.7},
$4=mr^{-}(\mathbb{F},G)\geq 2p+4m$. If $m=1$, then $p=0$, a
contradiction to the fact that $v$ is a cut vertex of $G$. Thus
$m=0$, implying that $p=2$. Let $W_{1}$, $W_{2}$ be the vertex sets
of the two complete multipartite  components of $G-v$ and let
$G_{1}$, $G_{2}$ be the subgraph induced by $\{v\}\cup W_{1}$,
$\{v\}\cup W_{2}$. By Lemma~\ref{lm2.5},
$mr^{-}(\mathbb{F},G_{1}-v)=mr^{-}(\mathbb{F},G_{2}-v)=2$. By
Lemma~\ref{lm2.7},
$4=mr^{-}(\mathbb{F},G)=mr^{-}(\mathbb{F},G_{1}-v)+mr^{-}(\mathbb{F},G_{2}-v)+\min\{r_{v}^{-}(\mathbb{F},G_{1})+r_{v}^{-}(\mathbb{F},G_{2}),2\}=2+2+\min\{r_{v}^{-}(\mathbb{F},G_{1})+r_{v}^{-}(\mathbb{F},G_{2}),2\}$.
Then $r_{v}^{-}(\mathbb{F},G_{1})=r_{v}^{-}(\mathbb{F},G_{2})=0$.
Thus $mr^{-}(\mathbb{F},G_{1})=mr^{-}(\mathbb{F},G_{2})=2$.  By
Lemma~\ref{lm2.5}, $G_1$ and $G_2$ are complete multipartite graphs,
and then (i) follows.

\noindent {\bf Case 2.} $q\neq0$. Note that
$r_v^{-}(\mathbb{F},K_{2})=2$. By Lemma~\ref{lm2.7},
$4=mr^{-}(\mathbb{F},G)\geq 2p+4m+2$. Then $m=0$ and $p=1$, and thus
(ii) follows.
\end{proof}

Now we consider the minimum skew rank of $k$-paths.  Note that a
$k$-path with at least $k+2$ vertices has at least two vertices of
degree $k$ and any two vertices of degree $k$ are not adjacent. The
following lemma follows directly from the definition of $k$-path.

\begin{Lemma} \label{base} Let $G$ be a $k$-path with at least $k+2$
vertices, and $v$ a vertex of $G$ with degree $k$. Then $G-v$ is
also a $k$-path.
\end{Lemma}

%\begin{proof} If $|V(G)|=k+2$, then it is obvious. Suppose that $|V(G)|\ge k+3$.
%Obviously, $G$ is a $k$-tree. Thus $G-v$ is a $k$-tree. Let $u$ be
%any neighbor of $v$ in $G$. Suppose that $d_G(u)\ge k+2$. Then
%$d_{G-v}(u)\ge k+1$. Note that $G-v$ has at least two vertices of
%degree $k$. Thus $G$ has at least three vertices of degree $k$, a
%contradiction. It follows that $v$ has at least one neighbor in $G$
%of degree $k+1$.  If $v$ has two neighbors of degree $k+1$ in $G$,
%then they become adjacent vertices of degree $k$ in $G-v$, which is
%impossible because $G-v$ is a $k$-tree. Thus $v$ has exactly one
%neighbor of degree $k+1$ in $G$. Then $G-v$ has exactly two vertices
%of degree $k$, and thus $G-v$ is a $k$-path.
%\end{proof}

Let $G$ be a $k$-path with $n\ge k+2$ vertices. By Lemma~\ref{base},
 the vertices of $G$ may be labeled as follows: Choose a vertex
of degree $k$, labeled as $v_n$, and label its unique neighbor of
degree $k+1$ in $G$ with $v_{n-1}$. Then $v_{n-1}$ is a vertex of
degree $k$ in the $k$-path  $G-v_n$. Repeating the process above, we
may label $n-{k+1}$ vertices of $G$ as $v_n$, $v_{n-1}, \dots,
v_{k+2}$. Obviously, $G-v_n-v_{n-1}-\cdots -v_{k+2}=K_{k+1}$ and it
contains a vertex of degree $k$ in $G$, which is labeled as $v_1$,
and the remaining vertices are labeled as $v_2$, $v_3, \dots,
v_{k+1}$ such that $v_2$ is the unique neighbor of $v_1$ with degree
$k+1$ in $G$. Note that in our labelling, $v_i$ is not adjacent to
$v_{j+1}, v_{j+2},\dots, v_n$ if $v_i$ is not adjacent to $v_j$ for
$j\ge \max\{i+1,k+2\}$.  Recall that a $k$-tree is a chordal graph. The above labeling is
actually the ``perfect elimination" labeling inherent to chordal graphs \cite{Sh}.% We
%use this labelling in the proof of the following theorem.

\begin{Theorem} \label{th1}
Let $G$ be a $k$-$path$ on $n$ vertices and  $\mathbb{F}$  an
infinite field. Then
\begin{equation*}
mr^{-}(\mathbb{F},G)=\begin{cases}n-k &\text{if $n-k$ is even},\\
                       n-k+1 &\text{if $n-k$ is odd}.
\end{cases}
\end{equation*}
\end{Theorem}

\begin{proof}
Let $Z=\{v_{1},v_{2},\ldots , v_{k}\}$. Color all vertices in $Z$
black and all the vertices outside $Z$ white. We will show that $Z$
is a zero forcing set of $G$. Since all neighbors of $v_{1}$
different from $v_{k+1}$ are black, we have $v_{1}\rightarrow
v_{k+1}$. Note that $v_{2}$ is adjacent to $v_{k+2}$ but not
adjacent to $v_{k+3},v_{k+4},\ldots, v_{n}$.
 Since all neighbors of $v_{2}$
different from $v_{k+2}$ are black, we have $v_{2}\rightarrow
v_{k+2}$. Let $G_1=G[\{v_{1},v_{2},\ldots , v_{k+3}\}]$ and
$G_2=G[\{v_{1},v_{2},\ldots , v_{k+4}\}]$. If each neighbor of
$v_{k+3}$ in $G_1$ is adjacent to $v_{k+4}$ in $G$, then $v_{k+4}$
is of degree $k+1$ in $G_2$, a contradiction. Thus there is a
neighbor, say $w$, of $v_{k+3}$ in $G_1$ such that $wv_{k+4}\not\in
E_G$, and then $wv_i\not\in E_G$ for $i\ge k+5$, implying that
 $w\rightarrow v_{k+3}$.
 Repeating the process above, we may finally color all vertices of $G$ black. Thus $Z$ is
a zero forcing set of $G$.  By Lemma~\ref{lm2.4},
$M^{-}(\mathbb{F},G)\leq Z(G)\leq k$, and then
$mr^{-}(\mathbb{F},G)=n-M^{-}(\mathbb{F},G)\geq n-k$. Note that the
rank of a skew-symmetric matrix is even. It follows that
\begin{eqnarray*}
mr^{-}(\mathbb{F},G)\geq\begin{cases}n-k &\text{if $n-k$ is even},\\
                         n-k+1 &\text{if $n-k$ is odd}.
\end{cases}
\end{eqnarray*}

To prove the result, we need only to show
\begin{equation}\label{ee}
mr^{-}(\mathbb{F},G)\leq\begin{cases}n-k &\text{if $n-k$ is even},\\
                         n-k+1 &\text{if $n-k$ is odd}.
\end{cases}
\end{equation}
We prove this by induction on $n$. If $n=k+1$, then $G=K_{k+1}$,
which is a complete multipartite graph, and thus by
Lemma~\ref{lm2.5}, $mr^{-}(\mathbb{F},G)=2=n-k+1$. If $n=k+2$, then
$G=K_{k+2}-e$ is also a complete multipartite graph, where $e\in
E_{K_{k+2}}$, and thus by Lemma  \ref{lm2.5},
$mr^{-}(\mathbb{F},G)=2=n-k$. Thus (\ref{ee}) is true for $n=k+1,
k+2$. Suppose that $n\geq k+3$ and for a $k$-path $H$ on $m$
vertices with $k+1\leq m\leq n-1$, we have
\begin{equation*}
mr^{-}(\mathbb{F},H)\leq\begin{cases}m-k &\text{if $m-k$ is even},\\
                         m-k+1 &\text{if $m-k$ is odd}.
\end{cases}
\end{equation*}
Let $G$ be a $k$-path on $n$ vertices. Let
$$G_{1}=G[\{v_{1},v_{2},\ldots ,v_{k+2}\}] \mbox{ and }
G_{2}=G[\{v_{3},v_{4},\ldots ,v_{n}\}].$$ Then $G_{1}$ is a $k$-path
on $k+2$ vertices, and $G_{2}$ is a $k$-path on $n-2$ vertices.
Obviously, $mr^{-}(\mathbb{F},G_{1})=2$, and by the induction
hypothesis,
\begin{equation*}
mr^{-}(\mathbb{F},G_{2})\leq\begin{cases}n-k-2 &\text{if $n-k-2$ is even},\\
                         n-k-1 &\text{if $n-k-2$ is odd},
\end{cases}
\end{equation*}
i.e.,
\begin{equation*}
mr^{-}(\mathbb{F},G_{2})\leq\begin{cases}n-k-2 &\text{if $n-k$ is even},\\
                         n-k-1 &\text{if $n-k$ is odd}.
\end{cases}
\end{equation*}
Note that  $G=G_{1}\cup  G_{2}$. By Lemma~\ref{lm2.1},
\begin{eqnarray*}
mr^{-}(\mathbb{F},G)&\le& mr^{-}(\mathbb{F},G_{1})+ mr^{-}(\mathbb{F},G_{2})\\
         &\leq& 2+\begin{cases}n-k-2 &\text{if $n-k$ is even}\\
                         n-k-1 &\text{if $n-k$ is odd}
                         \end{cases}\\
         &=&\begin{cases}n-k &\text{if $n-k$ is even},\\
                         n-k+1 &\text{if $n-k$ is odd}.
                         \end{cases}
\end{eqnarray*}
This proves (\ref{ee}).
\end{proof}

%Let $\mathbb{R}$ be the real field.

%Recall that the $k$-th power $G^{k}$ of a graph $G$ is the graph
%whose vertex set is $V_G$, two distinct vertices being adjacent in
%$G^{k}$ if and only if their distance in $G$ is at most $k$. Let
%$P_n=v_1v_2\dots v_n$ be the path on $n$ vertices.
Obviously, $P_n^k$ is a complete graph if $k\ge n$. Suppose that
$k\le n-1$. Obviously, $P_n^k[\{v_1,v_2, \dots, v_{k+1}\}]=K_{k+1}$,
and if $k\le n-2$, then for $j=2,3,\dots,n-k$,
$P_n^k[\{v_{j},v_{j+1}, \dots, v_{k+j-1}\}]=K_{k}$, and $v_{k+j}$ is
adjacent to $v_{j},v_{j+1}, \dots, v_{k+j-1}$. Thus $P_{n}^{k}$ is a
$k$-path. Now by Lemma~\ref{lm2.5} and Theorem
\ref{th1} we have the following result, which was proved in \cite{DKT} when $\mathbb{F}$ is the real field $\mathbb{R}$.

\begin{Corollary} Let $\mathbb{F}$  be an
infinite field. Then
 % and let $P_{n}^{k}$ be the $k$-th power of $P_{n}$.
\begin{eqnarray*}
mr^{-}(\mathbb{F},P_{n}^{k})=\begin{cases}n-k &\text{if $1\le k\le n-1$ and $n-k$ is even},\\
                         n-k+1 &\text{if $1\le k\le n-1$ and $n-k$ is
                         odd},\\
                         2 &\text{if $k\ge n$}.

\end{cases}
\end{eqnarray*}
\end{Corollary}

%\begin{proof} Let $P_n=v_1v_2\dots v_n$.
%Obviously, $P_n^k$ is a complete graph if $k\ge n$. Suppose that
%$k\le n-1$. Obviously, $P_n^k[\{v_1,v_2, \dots, v_{k+1}\}]=K_{k+1}$,
%and if $k\le n-2$, then for $j=2,3,\dots,n-k$,
%$P_n^k[\{v_{j},v_{j+1}, \dots, v_{k+j-1}\}]=K_{k}$, and $v_{k+j}$ is
%adjacent to $v_{j},v_{j+1}, \dots, v_{k+j-1}$. Thus $P_{n}^{k}$ is a
%$k$-path. Then the result follows from Lemma~\ref{lm2.5} and Theorem
%\ref{th1}.
%\end{proof}

Finally, we gave an observation.

\begin{Theorem} \label{add}
Let  $G$ be a connected  graph with no even cycles and $\mathbb{F}$
a field. Then
 $mr^{-}(\mathbb{F},G)=2match(G)=MR^{-}(\mathbb{F},G)$.
\end{Theorem}

\begin{proof}
By Lemma \ref{lm2.8}, $mr^{-}(\mathbb{F},G)\leq
MR^{-}(\mathbb{F},G)=2match(G)$. Let $M$ be a maximum matching of
$G$ and $\{v_{1},\cdots,v_{k}\}$ the vertices in $M$. Then $M$ is a
perfect matching of $H=G[\{v_{1},\cdots,v_{k}\}]$. This perfect
matching is unique. Otherwise, the graph induced by the vertices of
the symmetric difference of two (different) perfect matchings of $H$
consists of even cycles, which is impossible because $G$ contains no
even cycles. By Lemmas \ref{lm2.9} and  \ref{lm2.6},
$mr^{-}(\mathbb{F},G)\geq mr^{-}(\mathbb{F},H)=2match(G)$. The
result follows.
\end{proof}

Note that a tree has no (even) cycles. By previous theorem we have the following result.

\begin{Corollary} \cite{AB}
Let $G$ be a tree and  $\mathbb{F}$  a field. Then
$mr^{-}(\mathbb{F},G)=2match(G)$ $=MR^{-}(\mathbb{F},G)$.
\end{Corollary}

 Let $G$ be a connected unicyclic graph with a unique cycle $C$. If $C$ is odd, then by Theorem \ref{add},
 $mr^{-}(\mathbb{F},G)=2match(G)$. Recall that it was shown in \cite{LMD} that if $C$ is odd, then 
 $mr^{-}(\mathbb{R},G)=2match(G)$, and if $C$ is even, then
 $mr^{-}(\mathbb{R},G)=2match(G)$ or $2match(G)-2$.

\bigskip
{\bf Acknowledgment.} This work was supported by the National
Natural Science Foundation of China (No.~11071089)and the Guangdong
Provincial Natural Science Foundation of China (no.~S2011010005539).

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\end{document}